OpenStudy (anonymous):

use the subitution method to solve the system equation y=3x-4 x=9

OpenStudy (anonymous):

first collect the similar terms

OpenStudy (anonymous):

If x equals 9. then plug 9 into the equation. y=3(9)-4 y=27-4 y=23 Now you have Y. 23=3(9)-4 23=23 I'm pretty confident that this is how you do it. If not, then I'm sorry

OpenStudy (anonymous):

thank you c:

OpenStudy (anonymous):

No problem